OpenStudy (anonymous):

Solve. 6xy^3z • 3x^2yx^2

OpenStudy (anonymous):

you'll have to use this formula: \[\LARGE x^n\cdot x^m=x^{n+m}\] ... example: \[\LARGE ax^n\cdot bx^m=(a\cdot b)x^{n+m}\]

OpenStudy (anonymous):

can you do it ?

OpenStudy (anonymous):

It Sent This : \[\LARGE x^n\cdot x^m=x^{n+m}\] ... example: \[\LARGE ax^n\cdot bx^m=(a\cdot b)x^{n+m}\]

OpenStudy (anonymous):

you did what? ... O_O did you copied my post? or retyped it? if you copied , how did you do that ? O_O :A :)

OpenStudy (anonymous):

I Copied It Ctrl+C

OpenStudy (anonymous):

do you know to solve your problem? (To me it didn't work just Ctrl+C ) :S

OpenStudy (anonymous):

(Ya Just Ctrl then C) I Don't Think I Can Solve It

OpenStudy (anonymous):

\[\LARGE 6xy^3z\cdot 3x^2yx^2=18x^{2+2+1}y^{3+1}z\] how about now ? :)

OpenStudy (anonymous):

6xy^3z * 3x^2yx^2 = 18x^(2+2+1)y^(3+1)z ...??

OpenStudy (anonymous):

sum exponents ... 18x^(2+2+1)y^(3+1)z =18x^(5)y^(4)z

OpenStudy (anonymous):

18x^(5)y^(4)z =18x^(5)y^(4)z ?

OpenStudy (anonymous):

18x^(5)y^(4)z is the final answer..

OpenStudy (anonymous):

Thank You For Ur Wonderful Help

OpenStudy (anonymous):

My pleasure ...